# On $[1,2]$-Domination in Interval and Circle Graphs

### v1

@@ Abstract
A subset $S$ of vertices in a graph $G=(V,E)$ is Dominating Set if each vertex in $V(G)\setminus S$ is adjacent to at least one vertex in $S$. Chellali et al. in 2013, by restricting the number of neighbors in $S$ of a vertex outside $S$, introduced the concept of $[1,j]$-dominating set. A set $D\subseteq V$ of a graph $G=(V,E)$ is called $[1,j]$-Dominating Set of $G$ if every vertex not in $D$ has at least one neighbor and at most $j$ neighbors in $D$. The Minimum $[1,j]$-Domination problem is the problem of finding the minimum set $D$. Given a positive integer $k$ and a graph $G=(V,E)$, the $[1,j]$-Domination Decision problem is to decide whether $G$ has $[1,j]$-dominating set of cardinality at most $k$. A polynomial-time algorithm was obtained in split graphs for a constant $j$ in contrast to the classic Dominating Set problem which is NP-hard in split graphs. This result motivates us to investigate the effect of restriction $j$ on the complexity of $[1,j]$-domination problem on various classes of graphs. Although for $j\geq 3$, it has been proved that the minimum of classical domination is equal to minimum $[1,j]$-domination in interval graphs, the complexity of finding the minimum $[1,2]$-domination in interval graphs is still outstanding. In this paper, we propose a polynomial-time algorithm for computing a minimum $[1,2]$ on non-proper interval graphs by a dynamic programming technique. Next, on the negative side, we show that the minimum $[1,2]$-dominating set problem on circle graphs is $NP$-complete.
@@ 1.1 Definition and notation
A circle graph is a graph $G=(V,E)$ such that there is a one-to-one mapping between vertices in $V$ and a set $C$ of chords in a circle. Two vertices in $V$ are adjacent if and only if the corresponding chords in $C$ intersect. $C$ is called the chord intersection model for $G$. Equivalently, the vertices of a circle graph can be placed in the one-to-one correspondence the intervals such that two vertices are adjacent if and only if the corresponding intervals overlap, but neither contains the other. $I$ is called the interval representation of the circle graph. There exists a polynomial transformation between a representation of a circle graph, the set of its chords or its intervals representation. Moreover those are equivalent.
@@ 2 Algorithm for interval graphs
Another proof for the above lemma was shown in [5]. The authors also show that for every interval graph $G$, $\gamma(G)=\gamma_{[1,3]}(G)$ and, there exists graph $G$ such that $\gamma(G)\leq\gamma_{[1,2]}(G)$. but the complexity of finding $[1,2]$-dominating set in interval graphs remained open.
@@
In the rest of this section, we assume that $G=(V,E)$ is an interval graph of order $n$, and without lose of generality with a numbering $(1,2,\ldots,n)$ of vertices of $G$ satisfying the condition of Corollary 1.
@@
Let $j\in[1,i)$, $k\in[j,i)$ and $b=\min\{\mathtt{low}(k+1,i),j\}$. If $\mathtt{maxlow}(i)\leq k$, then $\gamma^{0}_{[1,2]}(j,i:k)=\min\{\gamma^{1}_{[1,2]}(b,k:k),\gamma^{1}_{[1,2]}(b,k:k,l):l\in[b,j)\}$, otherwise, $\gamma^{0}_{[1,2]}(j,i:k)$ is not defined.
@@
Let $D$ be an $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality such that $i\in D$, $i^{\prime}=\max D\cap[1,i)$, $i^{\prime\prime}=\max D\cap[1,i^{\prime})$ and $\mathtt{low}(i)<i^{\prime\prime}$. Then, $x\in D$ for all $x\in[\mathtt{low}(i),i^{\prime\prime})$.
@@
Let $b=\mathtt{low}(i)$, $c=\mathtt{low}(j)$, $d=\mathtt{low}(j+1,i-1)$, $d^{\prime}=\mathtt{low}(b,j-1)$, $d^{\prime\prime}=\mathtt{low}(b,i-1)$, $z=\max\{x\in(j,i):d=\mathtt{low}(x)\}$, and $k^{\prime}=\min\{k,b,d,d^{\prime}\}$.
@@
Now, we are ready to present our algorithm (Algorithm 2.1) for computing $\gamma_{[1,2]}(G)$, where $G$ is an interval graph. Note that if $12\in E$, then the case (iv.b) of Lemma 7 holds and so $\gamma^{1}_{[1,2]}(1,2:2)<\gamma^{1}_{[1,2]}(1,2:2,1)$. By considering Lemmas 2, 3, 5, 6, 7 and 8 we see that if we use $\gamma^{1}_{[1,2]}(1,2:2,1)$ for computing a value, then we also use $\gamma^{1}_{[1,2]}(1,2:2)$ for computing that value. Hence, in Line 12 we set $\gamma^{1}_{[1,2]}(1,2:2,1)$ to $\infty$. We also use this note in Line 20 for computing $\gamma^{1}_{[1,2]}(k,i:i,j)$.
@@
1 Compute a numbering $(1,2,\ldots,n)$ of vertices of $G$ satisfying the condition of Corollary 1.
2 Compute $\mathtt{low}(i)$ for all $i\in[1,n]$.
3 Compute $\mathtt{low}(j,i)$ for all $1\leq j<i\leq n$.
4 Compute $\mathtt{maxlow}(i)$ for all $i\in[1,n]$.
5 $\gamma^{0}_{[1,2]}(1)\leftarrow\infty$;
6 If $12\in E$, then $\gamma^{0}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(2)\leftarrow\infty$;
7 $\gamma^{0}_{[1,2]}(1,2)\leftarrow\infty$;
8 If $12\in E$, then $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow\infty$;
9 $\gamma^{1}_{[1,2]}(1)\leftarrow 1$;
10 If $12\in E$, then $\gamma^{1}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(2)\leftarrow 2$;
11 If $12\in E$, then $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow 1$, otherwise, $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow\infty$;
12 $\gamma^{1}_{[1,2]}(1,2:2,1)\leftarrow\infty$;
13 $\gamma^{11}_{[1,2]}(1,3:3,2,1)\leftarrow 3$;
14 for $i=3$ to $n$ do
15       for $1\leq j<i$ do
16             Compute $\gamma^{0}_{[1,2]}(j,i)$ using Lemma 2.
17             Compute $\gamma^{1}_{[1,2]}(j,i:i)$ using Lemma 6.
18
19      for $1\leq k\leq j<i$ do
20             Compute $\gamma^{0}_{[1,2]}(k,i:j)$ using Lemma 3.
21             Compute $\gamma^{1}_{[1,2]}(k,i:i,j)$ using Lemma 7.
22
23      for $1\leq l<k^{\prime}<j<i$ do
24             Compute $\gamma^{1}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 8.
25             Compute $\gamma^{11}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 9.
26
27      Compute $\gamma^{0}_{[1,2]}(i)=\gamma^{0}_{[1,2]}(i,i)$ using Lemma 2.
28       Compute $\gamma^{1}_{[1,2]}(i)$ using Lemma 5.
29
30return $\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$;
Algorithm 2.1 [1,2]-Domination-Interval-$(G)$
@@ 13 $\gamma^{11}_{[1,2]}(1,3:3,2,1)\leftarrow 3$;
Let $G=(V,E)$ be an interval graph of order $n$. Algorithm 2.1 computes $\gamma_{[1,2]}(G)$ in $O(n^{4})$ time.
@@
Let $(1,2,\ldots,n)$ be a numbering of vertices of $G$ computed in Line 1. This numbering of vertices of $G=(V,E)$ can be computed in $O(|V|+|E|)$ time [12]. We need $O(n^{3})$ time to compute $\mathtt{low}(i)$ for all $i\in[1,n]$, $\mathtt{low}(j,i)$ for all $1\leq j<i\leq n$ and $\mathtt{maxlow}(i)$ for all $i\in[1,n]$ in Lines 2-4. It is easy to compute the values $\gamma^{0}_{[1,2]}(1)$, $\gamma^{0}_{[1,2]}(2)$, $\gamma^{0}_{[1,2]}(1,2)$, $\gamma^{0}_{[1,2]}(1,2:1)$, $\gamma^{1}_{[1,2]}(1)$, $\gamma^{1}_{[1,2]}(2)$, $\gamma^{1}_{[1,2]}(1,2:2)$, $\gamma^{1}_{[1,2]}(1,2:2,1)$ and $\gamma^{11}_{[1,2]}(1,3:3,2,1)$ as seen in Lines 5-13 in $O(1)$ time. Let $i\in[3,n]$, $j\in[1,i)$, $k\in[1,j]$, $k^{\prime}\in[2,j)$ and $l\in[1,k^{\prime})$. By Lemmas 2, 3, 5, 6, 7, 8 and 9, compute $\gamma^{0}_{[1,2]}(i)$, $\gamma^{0}_{[1,2]}(j,i)$, $\gamma^{0}_{[1,2]}(k,i:j)$, $\gamma^{1}_{[1,2]}(i)$, $\gamma^{1}_{[1,2]}(j,i:i)$, $\gamma^{1}_{[1,2]}(k,i:i,j)$, $\gamma^{1}_{[1,2]}(k,i:i,j,k^{\prime})$ and $\gamma^{11}_{[1,2]}(k,i:i,j,k^{\prime})$ in Lines 15-25. We see that the running time of Lines 14-25 is $O(n^{4})$. Hence, the running time of Algorithm 2.1 is $O(n^{4})$. Clearly, $\gamma_{[1,2]}(G)=\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$. This completes the proof of the theorem. ∎
@@ 3 $[1,2]$-domination in Circle Graphs
    Figure 10: Intervals corresponding each cluster
@@
    For each literal in clause $C_{j}$, the interval between the endpoint of $t_{j}^{l}$ and the endpoint of the first chords of $a_{j}^{l}$ is called the truth interval, and the interval between the right point of the second chord of $p_{j}^{l}$ and the right point of $f_{j}^{l}$ is called false interval for literal $l$ of clause $j$.
@@ 4 Conclusion
Applying a constraint that determines the maximum number of times a vertex outside the dominating set, like $D$, is dominated by the elements inside $D$, makes the problem easier or harder than the original domination problem. In this paper, first we have designed a polynomial algorithm for $[1,2]$-dominating problem in non-proper interval graphs. Although it seems $O(n^{4})$ is huge, it is an affirmative answer to the open problem on the complexity class of $[1,2]$-Domination problem in non-proper interval graphs. Designing more efficient algorithm for this problem is an interesting open problem. Then, we have shown that the $[1,2]$-Dominating Set problem is $NP$-complete on circle graphs. It is thus natural to ask for which subclass of circle graphs, like distance-hereditary or k-polygon, the problem has polynomial-time algorithm. Meanwhile, it would be desirable to show the complexity of $[1,2]$-dominating problem on other classes of graphs marked with a question mark in Figure 3.

### v2

@@ Abstract
A subset $S$ of vertices in a graph $G=(V,E)$ is Dominating Set if each vertex in $V(G)\setminus S$ is adjacent to at least one vertex in $S$. Chellali et al. in 2013, by restricting the number of neighbors in $S$ of a vertex outside $S$, introduced the concept of $[1,j]$-dominating set. A set $D\subseteq V$ of a graph $G=(V,E)$ is called $[1,j]$-Dominating Set of $G$ if every vertex not in $D$ has at least one neighbor and at most $j$ neighbors in $D$. The Minimum $[1,j]$-Domination problem is the problem of finding the minimum set $D$. Given a positive integer $k$ and a graph $G=(V,E)$, the $[1,j]$-Domination Decision problem is to decide whether $G$ has $[1,j]$-dominating set of cardinality at most $k$. A polynomial-time algorithm was obtained in split graphs for a constant $j$ in contrast to the classic Dominating Set problem which is NP-hard in split graphs. This result motivates us to investigate the effect of restriction $j$ on the complexity of $[1,j]$-domination problem on various classes of graphs. Although for $j\geq 3$, it has been proved that the minimum of classical domination is equal to minimum $[1,j]$-domination in interval graphs, the complexity of finding the minimum $[1,2]$-domination in interval graphs is still outstanding. In this paper, we propose a polynomial-time algorithm for computing a minimum $[1,2]$ on interval graphs by a dynamic programming technique. Next, on the negative side, we show that the minimum $[1,2]$-dominating set problem on circle graphs is $NP$-complete.
@@ 1.1 Definition and notation
A circle graph is a graph $G=(V,E)$ such that there is a one-to-one mapping between vertices in $V$ and a set $C$ of chords in a circle. Two vertices in $V$ are adjacent if and only if the corresponding chords in $C$ intersect. $C$ is called the chord intersection model for $G$. Equivalently, the vertices of a circle graph can be placed in the one-to-one correspondence the intervals such that two vertices are adjacent if and only if the corresponding intervals overlap, but neither contains the other. $I$ is called the interval representation of the circle graph. There exists a polynomial transformation between a representation of a circle graph, the set of its chords or its intervals representation. Moreover those are equivalent.(See Figure 2)
@@ 2 Algorithm for interval graphs
Another proof for the above lemma was shown in [5]. The authors also show that for every interval graph $G$, $\gamma(G)=\gamma_{[1,3]}(G)$ and, there exists graph $G$ such that $\gamma(G)\leq\gamma_{[1,2]}(G)$, but the complexity of finding $[1,2]$-dominating set in interval graphs remained open.
@@
In the rest of this section, we assume that $G=(V,E)$ is an interval graph of order $n$, and without loss of generality with a numbering $(1,2,\ldots,n)$ of vertices of $G$ satisfying the condition of Corollary 1.
@@
Let $i\in[1,n]$, $j\in[1,i)$, $k\in[j,i)$ and $b=\min\{\mathtt{low}(k+1,i),j\}$. If $\mathtt{maxlow}(i)\leq k$, then $\gamma^{0}_{[1,2]}(j,i:k)=\min\{\gamma^{1}_{[1,2]}(b,k:k),\gamma^{1}_{[1,2]}(b,k:k,l):l\in[b,j)\}$, otherwise, $\gamma^{0}_{[1,2]}(j,i:k)$ is not defined.
@@
Let $D$ be an $[1,2]$-dominating set of $G[1,i]$ with minimum cardinality such that $i\in[1,n]$, $i\in D$, $i^{\prime}=\max D\cap[1,i)$, $i^{\prime\prime}=\max D\cap[1,i^{\prime})$ and $\mathtt{low}(i)<i^{\prime\prime}$. If both $i^{\prime}$ and $i^{\prime\prime}$ exist, then $x\in D$ for all $x\in[\mathtt{low}(i),i^{\prime\prime})$.
@@
Let $i\in[1,n]$, $j\in[1,i)$, $b=\mathtt{low}(i)$, $c=\mathtt{low}(j)$, $d=\mathtt{low}(j+1,i-1)$, $d^{\prime}=\mathtt{low}(b,j-1)$, $d^{\prime\prime}=\mathtt{low}(b,i-1)$, $z=\max\{x\in(j,i):d=\mathtt{low}(x)\}$, and $k^{\prime}=\min\{k,b,d,d^{\prime}\}$.
@@
Let $i\in[1,n]$, $j\in[1,i)$, $k\in[1,j]$, $b=\mathtt{low}(i)$, $c=\mathtt{low}(j)$, $d=\mathtt{low}(k)$, $e=\mathtt{low}(j+1,i-1)$, $e^{\prime}=\mathtt{low}(b,i-1)$ and $z=\max\{x\in(j,i):e=\mathtt{low}(x)\}$.
@@
Now, we are ready to present our algorithm (Algorithm 1) for computing $\gamma_{[1,2]}(G)$, where $G$ is an interval graph. Note that if $12\in E$, then the case (iv.b) of Lemma 7 holds and so $\gamma^{1}_{[1,2]}(1,2:2)<\gamma^{1}_{[1,2]}(1,2:2,1)$. By considering Lemmas 2, 3, 5, 6, 7 and 8 we see that if we use $\gamma^{1}_{[1,2]}(1,2:2,1)$ for computing a value, then we also use $\gamma^{1}_{[1,2]}(1,2:2)$ for computing that value. Hence, in Line 12 we set $\gamma^{1}_{[1,2]}(1,2:2,1)$ to $\infty$. We also use this note in Line 20 for computing $\gamma^{1}_{[1,2]}(k,i:i,j)$.
Algorithm 1 [1,2]-Domination-Interval-$(G)$
@@
1:Compute a numbering $(1,2,\ldots,n)$ of vertices of $G$ satisfying the condition of Corollary 1.
2:Compute $\mathtt{low}(i)$ for all $i\in[1,n]$.
3:Compute $\mathtt{low}(j,i)$ for all $1\leq j<i\leq n$.
4:Compute $\mathtt{maxlow}(i)$ for all $i\in[1,n]$.
5:$\gamma^{0}_{[1,2]}(1)\leftarrow\infty$
6:If $12\in E$, then $\gamma^{0}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(2)\leftarrow\infty$
7:$\gamma^{0}_{[1,2]}(1,2)\leftarrow\infty$
8:If $12\in E$, then $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow 1$, otherwise, $\gamma^{0}_{[1,2]}(1,2:1)\leftarrow\infty$
9:$\gamma^{1}_{[1,2]}(1)\leftarrow 1$
10:If $12\in E$, then $\gamma^{1}_{[1,2]}(2)\leftarrow 1$, otherwise, $\gamma^{1}_{[1,2]}(2)\leftarrow 2$
11:If $12\in E$, then $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow 1$, otherwise, $\gamma^{1}_{[1,2]}(1,2:2)\leftarrow\infty$
$\gamma^{1}_{[1,2]}(1,2:2,1)\leftarrow\infty$
$\gamma^{11}_{[1,2]}(1,3:3,2,1)\leftarrow 3$
14:for $i=3$ to $n$ do
15:     for $1\leq j<i$ do
16:         Compute $\gamma^{0}_{[1,2]}(j,i)$ using Lemma 2.
17:         Compute $\gamma^{1}_{[1,2]}(j,i:i)$ using Lemma 6.
18:     end for
19:     for $1\leq k\leq j<i$ do
20:         Compute $\gamma^{0}_{[1,2]}(k,i:j)$ using Lemma 3.
21:         Compute $\gamma^{1}_{[1,2]}(k,i:i,j)$ using Lemma 7.
22:     end for
23:     for $1\leq l<k^{\prime}<j<i$ do
24:         Compute $\gamma^{1}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 8.
25:         Compute $\gamma^{11}_{[1,2]}(l,i:i,j,k^{\prime})$ using Lemma 9.
26:     end for
27:     Compute $\gamma^{0}_{[1,2]}(i)=\gamma^{0}_{[1,2]}(i,i)$ using Lemma 2.
28:     Compute $\gamma^{1}_{[1,2]}(i)$ using Lemma 5.
29:end for
30:return $\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$
@@
Let $G=(V,E)$ be an interval graph of order $n$. Algorithm 1 computes $\gamma_{[1,2]}(G)$ in $O(n^{4})$ time.
@@
Let $(1,2,\ldots,n)$ be a numbering of vertices of $G$ computed in Line 1. This numbering of vertices of $G=(V,E)$ can be computed in $O(|V|+|E|)$ time [12]. We need $O(n^{3})$ time to compute $\mathtt{low}(i)$ for all $i\in[1,n]$, $\mathtt{low}(j,i)$ for all $1\leq j<i\leq n$ and $\mathtt{maxlow}(i)$ for all $i\in[1,n]$ in Lines 2-4. It is easy to compute the values $\gamma^{0}_{[1,2]}(1)$, $\gamma^{0}_{[1,2]}(2)$, $\gamma^{0}_{[1,2]}(1,2)$, $\gamma^{0}_{[1,2]}(1,2:1)$, $\gamma^{1}_{[1,2]}(1)$, $\gamma^{1}_{[1,2]}(2)$, $\gamma^{1}_{[1,2]}(1,2:2)$, $\gamma^{1}_{[1,2]}(1,2:2,1)$ and $\gamma^{11}_{[1,2]}(1,3:3,2,1)$ as seen in Lines 5-13 in $O(1)$ time. Let $i\in[3,n]$, $j\in[1,i)$, $k\in[1,j]$, $k^{\prime}\in[2,j)$ and $l\in[1,k^{\prime})$. By Lemmas 2, 3, 5, 6, 7, 8 and 9, compute $\gamma^{0}_{[1,2]}(i)$, $\gamma^{0}_{[1,2]}(j,i)$, $\gamma^{0}_{[1,2]}(k,i:j)$, $\gamma^{1}_{[1,2]}(i)$, $\gamma^{1}_{[1,2]}(j,i:i)$, $\gamma^{1}_{[1,2]}(k,i:i,j)$, $\gamma^{1}_{[1,2]}(k,i:i,j,k^{\prime})$ and $\gamma^{11}_{[1,2]}(k,i:i,j,k^{\prime})$ in Lines 15-29 We see that the running time of Lines 14-29 is $O(n^{4})$. Hence, the running time of Algorithm 1 is $O(n^{4})$. Clearly, $\gamma_{[1,2]}(G)=\min\{\gamma^{0}_{[1,2]}(n),\gamma^{1}_{[1,2]}(n)\}$. This completes the proof of the theorem. ∎
@@ 3 $[1,2]$-domination in Circle Graphs
    Figure 10: Intervals corresponding each clause
@@
    For each literal in clause $C_{j}$, the interval between the endpoint of $t_{j}^{l}$ and the endpoint of the first chords of $a_{j}^{l}$ is called the truth interval, and the interval between the right point of the second chord of $p_{j}^{l}$ and the right point of $f_{j}^{l}$ is called false interval for literal $l$ of clause $j$. (See Figure 11)
@@ 4 Conclusion
Applying a constraint that determines the maximum number of times a vertex outside the dominating set, like $D$, is dominated by the elements inside $D$, makes the problem easier or harder than the original domination problem. In this paper, first we have designed a polynomial algorithm for $[1,2]$-dominating problem in proper interval graphs. Although it seems $O(n^{4})$ is huge, it is an affirmative answer to the open problem on the complexity class of $[1,2]$-Domination problem in non-proper interval graphs. Designing more efficient algorithm for this problem is an interesting open problem. Then, we have shown that the $[1,2]$-Dominating Set problem is $NP$-complete on circle graphs. It is thus natural to ask for which subclass of circle graphs, like distance-hereditary or k-polygon, the problem has polynomial-time algorithm. Meanwhile, it would be desirable to show the complexity of $[1,2]$-dominating problem on other classes of graphs marked with a question mark in Figure 3.